dy/dx+2y/x=sinx/x²


1
Expert's answer
2021-05-17T15:27:34-0400
"x^2y'+2xy=\sin x"

"(x^2y)'=\sin x"

"x^2y=-\cos x+c_1"

"y=\dfrac{c_1}{x^2}-\dfrac{\cos x}{x^2}"